Carla Company, which is subject to a 40% income tax rate, projected its income before taxes for next year as shown here:

Sales (184,000 units) $9,200,000
Cost of sales
Variable costs 2,300,000
Fixed costs

3,450,000

Pretax earning

$3,450,000

If Carla wants after-tax earnings of 30% of sales, what is the required level of sales in dollars and in units?

Revenue: $

  

or

units
Answer #1

Desired after tax earnings = 30% of Sales

Before tax = 30% of Sales/(1-tax rate)

= 30% of Sales/(1-40%)

= 50% of Sales

Sales - variable costs - fixed costs = Pretax earnings

Sales - 25% of Sales - 3,450,000 = 50% of Sales

Sales = 3,450,000/25%

= $13,800,000

Hence, required revenue = $13,800,000

in Units = 13,800,000/50 = 276000 units